# Comparing Specialised Small and General Large Language Models on Text Classification: 100 Labelled Samples to Achieve Break-Even Performance

### v3

@@ 1 Introduction
So far, most of the studies comparing the performance of specialised and general language models from the data-efficiency perspective have been done using different (often non-representative and incomplete) settings and methodologies, leading to divergent findings (Pecher et al., 2024b). The specialised language models are tuned either on the whole labelled set (where they achieve higher performance than general language models) Schick and Schütze (2021); Qin et al. (2023), or using the same set of few samples (e.g., $4-32$) used for general models (where they achieve significantly worse performance than general models) Ma et al. (2023). Only few works study the behaviour of models between the few samples and the whole set, and observe the performance break-even points where specialised language models outperform general ones. Moreover, they often focus only on specific approaches or ignore the performance variance (caused by various sources of randomness, such as random seeds) Le Scao and Rush (2021); Hongjin et al. (2022); Hernandez et al. (2023); Gupta et al. (2023); Pecher et al. (2024b).
@@
Our goal is to remedy these shortcomings and answer the following question: "How many labelled samples do the specialised smaller models need to outperform more general larger models?" To achieve this, we investigate and observe how the performance of different approaches changes when increasing the number of labelled training samples (data-efficiency). We identify the break-even points between performance of specialised models and more general models of different sizes (compute-efficiency), while taking the performance variance into consideration (sensitivity to randomness). The aggregated results are presented in Figure 1. Our main contributions and findings are [^1] To support replicability and extension of our results, we openly publish the source code of our experiments at https://anonymous.4open.science/r/investigation-with-interactions-F6D4:
@@
  • Based on further analysis we find following key insights: 1) the required number of labelled samples is dependent on the dataset characteristics, with binary datasets and datasets that require better language understanding (e.g., question answering) requiring more labelled samples; 2) larger models do not necessarily lead to better model performance, data-efficiency or lower performance variance for general models; and 3) 4-bit quantisation has negligible impact on the overall behaviour of general models
@@ 2 Related Work
If enough labelled samples are used, the smaller specialised models can achieve performance on par with, or in many cases even better than the performance of the larger general models Schick and Schütze (2021); Qin et al. (2023); Hernandez et al. (2023). At the same time, in the extremely limited settings, where the models are fine-tuned using the same low number of samples as it is used for in-context learning, the general large language models excel over the small specialised models Ma et al. (2023); Hongjin et al. (2022). Other papers study the impact of varying the training data sizes on the performance of specialised models and their comparison with general models. However, they often either perform a comprehensive comparison across multiple approaches, but focus only on few samples and vary the data sizes only to up 32 samples Logan IV et al. (2022); Gao et al. (2021), or focus on small subset of approaches while using a larger part or the whole dataset, but often in specific domains Le Scao and Rush (2021); Gupta et al. (2023); Hongjin et al. (2022); Hernandez et al. (2023); Ma et al. (2023). Only few papers consider (to a certain extent) effects of different systematic choices (e.g., number of samples, or prompt format) or sensitivity to sources of randomness (e.g., choice or order of samples) on comparison (Ma et al., 2023; Pecher et al., 2024a; Sclar et al., 2023; Weber et al., 2023).
@@ 3 Comparison Methodology: Identifying Performance Break-Even Points
To further explore the impact of the sensitivity to the effects of randomness on the comparison, we identify the second break-even points (average+variance). This break-even point denotes the "worst-case size" or the point after which the specialised models show better performance even when the variance is taken into consideration. In other words, there is low probability that the performance of the specialised models will be lower even when using low number of runs or cherry-picking them. To identify this break-even point, we repeat the training and evaluation multiple times and then compare the worst case performance of the better performing model (obtained by subtracting the standard deviation from the mean) with the best case performance of the worse performing model (obtained by adding the standard deviation to the mean). Each repeat uses the same fixed random seed for all models to guarantee deterministic and replicable results. The random seed covers all sources of randomness (Gundersen et al., 2022a; Pecher et al., 2024b) – the split of data, selection of labelled samples, initialisation of models, order of data in training, non-deterministic operations (e.g., dropout), and selection of samples for in-context learning.
@@
We investigate the impact of training dataset size across a set of currently popular approaches for dealing with limited labelled data in NLP: 1) fine-tuning; 2) instruction-tuning; 3) prompting; and 4) in-context learning. For fine-tuning, we use the BERT Devlin et al. (2019a) and RoBERTa Liu et al. (2019) base models. For prompting and in-context learning the Flan-T5 Chung et al. (2024) base, GPT (4o-mini-2024-07-18 version) Brown et al. (2020); Ouyang et al. (2022), LLaMA-2 Touvron et al. (2023) 13B chat optimised model, LLaMA-3 (Dubey et al., 2024) 8B instruction optimised model, Mistral-7B Jiang et al. (2023) and Zephyr-7B Tunstall et al. (2023) models. For instruction-tuning we use the Flan-T5 (full instruction-tuning), Mistral-7B and Zephyr-7B models (both instruction-tuned using LoRA Hu et al. (2021)). Due to the significant costs, we use 4-bit quantisation for the LLaMA-2 model. In case of Mistral-7B and Zephyr-7B, we use both the full-precision and the 4-bit quantised versions, reporting only the best performance.
@@
Each experiment is repeated 100 times for fine-tuning and 20 times for the remaining approaches (due to the significant costs of inference or training of the larger language models). For in-context learning, we use as many samples per class as the context-size of the models allow. As increasing this number after a certain point results in degraded performance, we always report the best performance achieved up until the dataset subset. For further details see Appendix A
@@ 3.1 Comparison of Specialised and General Models
Figure 2: The impact of varying size of available labelled samples (in logarithmic scale) on the performance of fine-tuning, prompting, in-context learning and instruction-tuning approaches, reported using F1 macro and its standard deviation. For each approach, we select only the best performing model. We can observe that specialised models can often outperform general models with only a relatively small number of labelled samples ($10-1000$).
@@
Specialised models can outperform the general ones using only a relatively small number of labelled examples. To outperform zero-shot setting (prompting), fine-tuning approaches require on average between $10-500$ labelled samples. In case of few-shot setting (in-context learning), the number of required labelled samples is higher, on average between $100-2000$ samples.
@@
However, the comparison between models and the number of required labelled samples is strongly affected by the dataset and task characteristics. This includes characteristics like 1) how many classes are used (binary vs. multi-class setting); 2) length of sentences (e.g., SST2 vs. CoLA); 3) whether the task requires working with a single sentence or a paired/multiple inputs (e.g., SST2/CoLA vs. MRPC/BoolQ); or 4) the overall type of the task (e.g., simple sentiment classification vs. question answering) that is defined by these characteristics. The largest difference can be observed between binary and multi-class datasets. On multi-class datasets, fine-tuning requires only up to $100-200$ labelled samples to outperform the larger models (even instruction-tuned ones in many cases). On binary datasets the required samples are as high as $4000$ to outperform zero-shot setting (prompting), up to $9000$ to outperform few-shot setting (in-context learning), or up to $20000$ to outperform the instruction-tuned models. The remaining characteristics have lower impact that is mostly dependent on the model size (more information is provided in Appendix B).
@@ 3.2 Impact of Performance Variance on Comparison
CoLA $350~{}|~{}700^{+100\%}$ $500~{}|~{}NA$ $1800~{}|~{}NA$
BoolQ $4250~{}|~{}6000^{+41\%}$ $5000~{}|~{}NA$ $5000~{}|~{}7500^{+50\%}$
AGNews $40~{}|~{}60^{+50\%}$ $300~{}|~{}800^{+167\%}$ $2500~{}|~{}NA$
@@
Sensitivity to the effects of randomness, and the performance variance it causes, has a significant effect on the break-even points between models, increasing the number of required labelled samples by a significant margin. On average, fine-tuning requires $2-3$ times more labelled samples (increase of $100-200\%$) to outperform the remaining approaches when taking the variance into consideration. In specific cases, the impact of variance is negligible, as we observe a $0\%$ increase in the number of required samples (e.g., prompting on DB Pedia dataset). At the same time, the impact of variance is significantly higher in other cases, leading to an increase of up to $500\%$, or an increase where even with the full labelled dataset the second break-even point is not achieved (e.g., in-context learning on CoLA, BoolQ or SNIPS datasets).
The impact of the performance variance strongly depends on the dataset and the overall number of labelled samples required for the first break-even point. Looking at the increase in absolute numbers, an increase of $100-200\%$ represents the need to annotate only between $10-600$ more samples in some cases (e.g., most multi-class datasets), while representing an increase of $2000-11000$ or more labelled samples in other cases (e.g., binary datasets, mainly SST2). However, the impact is most significant in cases where the second break-even point does not exist, but can be obscured. For example on the SNIPS datasets, the first break-even point with in-context learning is achieved at $50$ labelled samples, but the second is not observed even when using all the $15000$ labelled samples of the dataset (corresponding to an increase larger than $3000\%$).
@@ 3.3 Additional Analysis: Effects of Model Size and Quantisation
Larger models do not consistently lead to better performance or lower performance variance. The impact of model size is most explicit for fine-tuning, considerably increasing the performance and reducing variance. For prompting and in-context learning, we often observe smaller models outperforming their larger counterparts in both the performance and its variance. At the same time, larger general models often benefit more from using larger number of in-context examples. However, this is mainly effect of their larger context size (i.e., can handle more samples before observing drop in performance) and not size of the model. For more details see Appendix B.
Using 4-bit quantisation has minimal impact on the overall sensitivity and performance of the general models. The quantised versions of the models often outperform their full-precision counterparts when used for in-context learning. On the other hand, for prompting, the full-precision models perform slightly better. Finally, 4-bit quantised models show the same level of performance variance and benefit the same from using more in-context examples as the full-precision models. For more details see Appendix C.
@@ 4 Discussion: Recommendations Based on the Findings
Instruction-tuning of the larger language models is the optimal solution as it provides consistent benefits for all the other cases (large computation budget with either small or large annotation budget). Instruction tuning provides the best trade-off for the performance and the required resources – increasing the number of labelled samples does not have a significant impact on the performance, while increasing the size of the instruction-tuned model has only low performance impact. However, if interested in best overall performance on the task, using the largest general model and using as many labelled samples for instruction-tuning (and in-context examples) should provide the most benefit even at the cost of significant increase in effort (and strong diminishing returns). In addition, this approach requires significantly larger computation budget than fine-tuning, due to the inference costs, which should be considered in the long run.
@@
Finally, when using general models, a specific focus should be on the prompt format and the set of in-context examples to maximise their benefits. The prompt format was identified as the most significant contributor to the performance and variance in the results. To achieve fair comparison and robustness of the models, the comparison should be done using prompt optimised for each model or an ensemble of multiple prompts for instruction tuning and evaluation. In addition, choosing the optimal number of in-context examples is important to achieve fair comparison. Using all the available samples may not always be possible or beneficial due to the limited context size. As illustrated in our experiments, the performance drops significantly after this limit is reached (or sometimes even beforehand). Furthermore, the quality of the samples is as important (if not more) than the quantity of the samples. Majority of the observed variance in the performance in our experiments is the result of the random choice of examples. Previous studies have also observed that the quality of the samples is paramount, even though identifying the samples of "highest quality" may not be as straightforward (Pecher et al., 2024c; Agrawal et al., 2023; Chang and Jia, 2023). As such, both quality and quantity of the in-context examples should be considered in the comparisons, introducing further human labour that should be considered when choosing the most effective and efficient approach and model.
@@ 5 Conclusion
In this paper, we perform a comprehensive and fair comparison of currently popular approaches for data-efficient learning in NLP, from the perspective of their data-efficiency and sensitivity to the effects of randomness. The main focus of the investigation is to determine how many labelled training samples are needed for the specialised small language models (obtained through fine-tuning or instruction-tuning) to outperform their general larger counterparts used in zero and few-shot settings. Based on the break-even points from 8 representative text classification datasets of various characteristics, we find that the number of required training samples is quite small, but strongly dependent on the dataset and task characteristics, with more labelled samples needed for binary datasets and tasks that require better language understanding. In addition, we identify a significant influence of the performance variance, stemming from the sensitivity of the different approaches, especially in-context learning and fine-tuning with few labelled samples, on the overall comparison and the number of required labelled samples. We can conclude that the specialised smaller models are a strong contender on the text classification tasks, with the general language models showing their benefits for specific cases, such as quick prototyping or when faced with extremely limited annotation or computation budget. Finally, based on findings of our experiments, we provide recommendations that allow for better decision making regarding what approach to use for different settings.

### v4

@@ 1 Introduction
So far, most of the studies comparing the performance of specialised and general language models from the data-efficiency perspective have been done using different (often non-representative and incomplete) settings and methodologies, leading to divergent findings (Pecher et al., 2024b). The specialised language models are tuned either on the whole labelled set (where they achieve higher performance than general language models) Schick and Schütze (2021); Qin et al. (2023), or using the same set of few samples (e.g., 4324-32) used for general models (where they achieve significantly worse performance than general models) Ma et al. (2023). Only few works study the behaviour of models between the few samples and the whole set, and observe the performance break-even points where specialised language models outperform general ones. Moreover, they often focus only on specific approaches or ignore the performance variance (caused by various sources of randomness, such as random seeds) Le Scao and Rush (2021); Hongjin et al. (2022); Hernandez et al. (2023); Gupta et al. (2023); Pecher et al. (2024b).
@@
Our goal is to remedy these shortcomings and answer the following question: "How many labelled samples do the specialised smaller models need to outperform more general larger models?" To achieve this, we investigate and observe how the performance of different approaches changes when increasing the number of labelled training samples (data-efficiency). We identify the break-even points between performance of specialised models and more general models of different sizes (compute-efficiency), while taking the performance variance into consideration (sensitivity to randomness). The aggregated results are presented in Figure 1. Our main contributions and findings are [^1] To support replicability and extension of our results, we openly publish the source code of our experiments at https://github.com/kinit-sk/L3D-sensitivity-investigation:
@@
  • Based on further analysis we find following key insights: 1) the required number of labelled samples is dependent on the dataset characteristics, with binary datasets and datasets that require better language understanding (e.g., question answering) requiring more labelled samples; 2) larger models do not necessarily lead to better model performance, data-efficiency or lower performance variance for general models; and 3) 4-bit quantisation has negligible impact on the overall behaviour of general models.
@@ 2 Related Work
If enough labelled samples are used, the smaller specialised models can achieve performance on par with, or in many cases even better than the performance of the larger general models Schick and Schütze (2021); Qin et al. (2023); Hernandez et al. (2023). At the same time, in the extremely limited settings, where the models are fine-tuned using the same low number of samples as it is used for in-context learning, the general large language models excel over the small specialised models Ma et al. (2023); Hongjin et al. (2022). Other papers study the impact of varying the training data sizes on the performance of specialised models and their comparison with general models. However, they often either perform a comprehensive comparison across multiple approaches, but focus only on few samples and vary the data sizes only to up 32 samples Logan IV et al. (2022); Gao et al. (2021), or focus on small subset of approaches while using a larger part or the whole dataset, but often in specific domains Le Scao and Rush (2021); Gupta et al. (2023); Hongjin et al. (2022); Hernandez et al. (2023); Ma et al. (2023). Only few papers consider (to a certain extent) effects of different systematic choices (e.g., number of samples, or prompt format) or sensitivity to sources of randomness (e.g., choice or order of samples) on comparison (Ma et al., 2023; Pecher et al., 2024c; Sclar et al., 2023; Weber et al., 2023).
@@ 3 Comparison Methodology: Identifying Performance Break-Even Points
To further explore the impact of the sensitivity to the effects of randomness on the comparison, we identify the second break-even points (average+variance). This break-even point denotes the "worst-case size" or the point after which the specialised models show better performance even when the variance is taken into consideration. In other words, there is low probability that the performance of the specialised models will be lower even when using low number of runs or cherry-picking them. To identify this break-even point, we repeat the training and evaluation multiple times and then compare the worst case performance of the better performing model (obtained by subtracting the standard deviation from the mean) with the best case performance of the worse performing model (obtained by adding the standard deviation to the mean). Each repeat uses the same fixed random seed for all models to guarantee deterministic and replicable results. The random seed covers all sources of randomness (Gundersen et al., 2022b; Pecher et al., 2024b) – the split of data, selection of labelled samples, initialisation of models, order of data in training, non-deterministic operations (e.g., dropout), and selection of samples for in-context learning.
@@
We investigate the impact of training dataset size across a set of currently popular approaches for dealing with limited labelled data in NLP: 1) fine-tuning; 2) instruction-tuning; 3) prompting; and 4) in-context learning. For fine-tuning, we use the BERT Devlin et al. (2019b) and RoBERTa Liu et al. (2019) base models. For prompting and in-context learning the Flan-T5 Chung et al. (2024) base, GPT-4 (4o-mini-2024-07-18 version) Brown et al. (2020); Ouyang et al. (2022), LLaMA-2 Touvron et al. (2023) 13B chat optimised model, LLaMA-3 (Dubey et al., 2024) 8B instruction optimised model, Mistral-7B Jiang et al. (2023) and Zephyr-7B Tunstall et al. (2023) models. For instruction-tuning we use the Flan-T5 (full instruction-tuning), Mistral-7B and Zephyr-7B models (both instruction-tuned using LoRA Hu et al. (2021)). Due to the significant costs, we use 4-bit quantisation for the LLaMA-2 model. In case of Mistral-7B and Zephyr-7B, we use both the full-precision and the 4-bit quantised versions, reporting only the best performance.
@@
Each experiment is repeated 100 times for fine-tuning and 20 times for the remaining approaches (due to the significant costs of inference or training of the larger language models). For in-context learning, we use as many samples per class as the context-size of the models allow. As increasing this number after a certain point results in degraded performance, we always report the best performance achieved up until the dataset subset. For further details see Appendix A.
@@ 3.1 Comparison of Specialised and General Models
Figure 2: The impact of varying size of available labelled samples (in logarithmic scale) on the performance of fine-tuning, prompting, in-context learning and instruction-tuning approaches, reported using F1 macro and its standard deviation. For each approach, we select only the best performing model. We can observe that specialised models can often outperform general models with only a relatively small number of labelled samples (10100010-1000).
@@
Specialised models can outperform the general ones using only a relatively small number of labelled examples. To outperform zero-shot setting (prompting), fine-tuning approaches require on average between 1050010-500 labelled samples. In case of few-shot setting (in-context learning), the number of required labelled samples is higher, on average between 1002000100-2000 samples.
@@
However, the comparison between models and the number of required labelled samples is strongly affected by the dataset and task characteristics. This includes characteristics like 1) how many classes are used (binary vs. multi-class setting); 2) length of sentences (e.g., SST2 vs. CoLA); 3) whether the task requires working with a single sentence or a paired/multiple inputs (e.g., SST2/CoLA vs. MRPC/BoolQ); or 4) the overall type of the task (e.g., simple sentiment classification vs. question answering) that is defined by these characteristics. The largest difference can be observed between binary and multi-class datasets. On multi-class datasets, fine-tuning requires only up to 100200100-200 labelled samples to outperform the larger models (even instruction-tuned ones in many cases). On binary datasets the required samples are as high as $4000$ to outperform zero-shot setting (prompting), up to $9000$ to outperform few-shot setting (in-context learning), or up to $20000$ to outperform the instruction-tuned models. The remaining characteristics have lower impact that is mostly dependent on the model size (more information is provided in Appendix B).
@@ 3.2 Impact of Performance Variance on Comparison
CoLA $350~|~700^{+100\%}$ 500|NA500~|~NA 1800|NA1800~|~NA
BoolQ $4250~|~6000^{+41\%}$ 5000|NA5000~|~NA $5000~|~7500^{+50\%}$
AGNews $40~|~60^{+50\%}$ $300~|~800^{+167\%}$ 2500|NA2500~|~NA
@@
Sensitivity to the effects of randomness, and the performance variance it causes, has a significant effect on the break-even points between models, increasing the number of required labelled samples by a significant margin. On average, fine-tuning requires 232-3 times more labelled samples (increase of $100-200\%$) to outperform the remaining approaches when taking the variance into consideration. In specific cases, the impact of variance is negligible, as we observe a $0\%$ increase in the number of required samples (e.g., prompting on DB Pedia dataset). At the same time, the impact of variance is significantly higher in other cases, leading to an increase of up to $500\%$, or an increase where even with the full labelled dataset the second break-even point is not achieved (e.g., in-context learning on CoLA, BoolQ or SNIPS datasets).
The impact of the performance variance strongly depends on the dataset and the overall number of labelled samples required for the first break-even point. Looking at the increase in absolute numbers, an increase of $100-200\%$ represents the need to annotate only between 1060010-600 more samples in some cases (e.g., most multi-class datasets), while representing an increase of 2000110002000-11000 or more labelled samples in other cases (e.g., binary datasets, mainly SST2). However, the impact is most significant in cases where the second break-even point does not exist, but can be obscured. For example on the SNIPS datasets, the first break-even point with in-context learning is achieved at $50$ labelled samples, but the second is not observed even when using all the $15000$ labelled samples of the dataset (corresponding to an increase larger than $3000\%$).
@@ 3.3 Additional Analysis: Effects of Model Size and Quantisation
Larger models do not consistently lead to better performance or lower performance variance. The impact of model size is most explicit for fine-tuning, considerably increasing the performance and reducing variance. For prompting and in-context learning, we often observe smaller models outperforming their larger counterparts in both the performance and its variance. At the same time, larger general models often benefit more from using larger number of in-context examples. However, this is mainly effect of their larger context size (i.e., can handle more samples before observing drop in performance) and not size of the model. In addition, for prompting and in-context learning, we do not observe that using larger models leads to lower performance variance. For more details see Appendix B.
Using 4-bit quantisation has minimal impact on the overall sensitivity and performance of the general models. The quantised versions of the models often outperform their full-precision counterparts when used for in-context learning (as can be seen in Figure 4). On the other hand, for prompting, the full-precision models perform slightly better, even though the difference in performance is minimal in almost all cases. Finally, 4-bit quantised models show the same level of performance variance and benefit the same from using more in-context examples as the full-precision models. In addition, we observe negligible impact of quantisation on the performance variance. Finally, quantisation has not impact on how much the models benefit from increasing the number of in-context examples or their context length. Therefore, we can conclude that the use of quantised models has minimal impact on the comparison between different approaches. As such, considering compute-efficiency and performance trade-off in the comparison, the quantised models should be preferred for the general large language models whenever possible. For more details, see Appendix C.
@@ 4 Discussion: Recommendations Based on the Findings
Instruction-tuning of the larger language models is the optimal solution as it provides consistent benefits for all the other cases (large computation budget with either small or large annotation budget). Instruction tuning provides the best trade-off for the performance and the required resources – increasing the number of labelled samples does not have a significant impact on the performance, while increasing the size of the instruction-tuned model has only low performance impact. However, if interested in best overall performance on the task, using the largest general model and using as many labelled samples for instruction-tuning (and in-context examples) should provide the most benefit at the cost of increase in effort (and strong diminishing returns). In addition, it requires significantly larger computation budget than fine-tuning, due to the inference costs, which should be considered in the long run.
@@
Finally, when using general models, a specific focus should be on the prompt format and the set of in-context examples to maximise their benefits. The prompt format was identified as the most significant contributor to the performance and variance in the results. To achieve fair comparison and robustness of the models, the comparison should be done using prompt optimised for each model or an ensemble of multiple prompts for instruction tuning and evaluation. In addition, choosing the optimal number of in-context examples is important to achieve fair comparison. Using all the available samples may not always be possible or beneficial due to the limited context size. As illustrated in our experiments, the performance drops significantly after this limit is reached (or sometimes even beforehand). Furthermore, the quality of the samples is as important (if not more) than the quantity of the samples. Majority of the observed variance in the performance in our experiments is the result of the random choice of examples. Previous studies have also observed that the quality of the samples is paramount, even though identifying the samples of "highest quality" may not be as straightforward (Pecher et al., 2024a; Agrawal et al., 2023; Chang and Jia, 2023). As such, both quality and quantity of the in-context examples should be considered in the comparisons, introducing further human labour that should be considered when choosing the most effective and efficient approach and model.
@@ 5 Conclusion
In this paper, we perform a comprehensive and fair comparison of currently popular approaches for data-efficient learning in NLP, from the perspective of their data-efficiency and sensitivity to the effects of randomness. The main focus of the investigation is to determine how many labelled training samples are needed for the specialised small language models (obtained through fine-tuning or instruction-tuning) to outperform their general larger counterparts used in zero and few-shot settings. Based on the break-even points from 8 representative text classification datasets of various characteristics, we find that the number of required training samples is quite small, but strongly dependent on the dataset and task characteristics, with more labelled samples needed for binary datasets and tasks that require better language understanding. In addition, we identify a significant influence of the performance variance, stemming from the sensitivity of the different approaches, especially in-context learning and fine-tuning with few labelled samples, on the overall comparison and the number of required labelled samples. We can conclude that the specialised smaller models are a strong contender on the text classification tasks, with the general language models showing their benefits for specific cases, such as quick prototyping or when faced with extremely limited annotation. Finally, based on our findings, we provide recommendations that allow for better decision making regarding what approach to use for different settings.